Southbound Company reported pretax net income from continuing operations of $800,000 and taxable income of $500,000. The book–tax difference of $300,000 was due to a $200,000 favorable temporary difference relating to depreciation, an unfavorable temporary difference of $80,000 due to an increase in the reserve for bad debts, and a $180,000 favorable permanent difference from the receipt of life insurance proceeds.

c. Compute Southbound Company’s effective tax rate.
Answer #1

Solution:-

Note:- Assuming tax rate is 34%

Effective tax rate:- 26.35%

Explanation:-

Pre-tax net income 800,000
Favorable temporary difference (200,000)
Unfavorable temporary difference 80,000
Favorable permanent difference (180,000)
Taxable income 500,000
Current income tax payable @34% 170,000
Favorable temporary difference (200,000)
Unfavorable temporary differences 80,000
Net favorable temporary difference (120,000)
Deferred tax expense (liability) @34% (40,800)

Total income tax provision = 170,000 + 40,800 = 210,800

Effective tax rate = 210,800 / 800,000 = 26.35%

Pelase Rate.